### v1

# Design and synthesis of scalable quantum programs

Abstract
We present a scalable, robust approach to creating quantum programs of arbitrary size and complexity. The approach is based on the true abstraction of the problem. The quantum program is expressed in terms of a high-level model together with constraints and objectives on the final program. Advanced synthesis algorithms transform the model into a low-level quantum program that meets the user’s specification and is directed at a stipulated hardware. This separation of description from implementation is essential for scale. The technology adapts electronic design automation methods to quantum computing, finding feasible implementations in a virtually unlimited functional space. The results show clear superiority over the compilation and transpilation methods used today. We expect that this technological approach will take over and prevail as quantum software become more demanding, complex, and essential.
Quantum algorithms are algorithms designed to work on digital quantum computers. Most modalities of quantum computers comply with a set of logical, universal gates, from which any algorithm can be assembled [1]. Implementing a specific quantum algorithm on a specific quantum hardware, requires maturity of the quantum software stack, including design languages, compilers, and technology to map the logical description into a concrete, efficient implementation. As quantum computers grow in power and resources, the problem of efficiently mapping logical algorithms to concrete implementations becomes increasingly difficult and must advance from the manual methods existing today [2]. In this paper, we present a new approach, radically different from regular compilation methods [3, 4, 5], to designing and implementing quantum algorithms. The approach, based on well-proven electronic design automation (EDA) principles and methods [6], is shown to reduce required resources (e.g. qubit count, 2-qubit gate count) by orders of magnitude compared to other existing approaches. This is done by automatically managing resources such as available qubits, multiple function implementations, functional levels of accuracy, and other high-level logic equivalences. Moreover, and in contrast to all existing approaches today, this approach is shown to be scalable to large-scale, realistic algorithms. Following the path of classical design automation, we anticipate that a significant design gap will follow the increasing power of quantum computers. Building powerful and scalable software technologies for quantum algorithm design is an essential step towards any meaningful quantum advantage. This work presents the most advanced steps taken in this direction to date, and lays the foundation for the future of the quantum software stack.
1 Introduction
Since 2017, we have seen massive and continuous advances in quantum hardware in many modalities, including superconducting devices [7, 8, 9, 10], ion traps [11, 12, 13], neutral atoms [14, 15], silicon devices [16, 17], photonic devices [18, 19], and more [20]. The consensus is that quantum computing has shifted from a scientific/academic phase into an engineering phase during this time frame. The different hardware roadmaps vary, but all show major applicative breakthroughs that are expected in the next few years [21, 22, 23, 24, 25, 26, 27, 28]. Software concepts and platforms have not evolved at the same pace. The first software platforms [4, 29, 30, 31] appeared hand-in-hand with the public exposition of quantum hardware accessible on the cloud in 2016 [32]. These platforms typically consisted of a means to specify the operations of quantum gates (and later quantum pulses) on a set of qubits, transpilation of the gate-level programs into equivalent gate-level programs suitable to a given hardware, and software for simulating the quantum hardware on classical CPUs and GPUs. Advancements in those software offerings included high-level constructs such as quantum registers and algorithmic building blocks. However, those were not true abstractions of the problem because the constructs only encapsulated hard-coded, low-level implementations of qubits and quantum gates; these were packaged as named-functions performing operations on named registers. Further developments in quantum software included application libraries built on top of the basic software, including applications in quantum machine learning, finance, chemistry, and many more [4, 5, 33, 34]. These libraries provided further domain-specific algorithms and functions, based on building blocks tuned and optimized for a specific application. Fundamentally, these were not different from hard-coded quantum circuits encapsulated in several levels of building block functions. In addition, several quantum programming languages that describe functionality at a higher level of abstraction in different respects have been proposed in recent years [3, 35, 36]. The tool chains for these languages utilize compilation frameworks and techniques adapted from classical programming language processing, notably MLIR [37] and LLVM [38]. The actual compilation approach in all these cases still rigidly maps the abstractions onto a low-level implementation, taking only local resource considerations into account. We present a radically different approach to the compilation of quantum programs. Our strategy follows and extends the general ideas presented earlier [39, 40, 2]. However, for the first time the ideas are fully implemented and do not remain at the conceptual levels. The level of understanding and detail gained through implementation clearly changes and cements the conceptual level as well; hence the work presented here creates a major step forward in advancing the path to true, scalable quantum software. At the base of our work is abstraction. By abstraction we mean allowing the user to express any information of interest to them regarding the quantum program, and omit any other information, or detail, that is not of interest. It is up to the technology to fill in all details that were not expressly specified, and come up with an executable, optimal, quantum program that conforms to the specification of the user and fits a given hardware. The abstract description from the user is a model of the program, while the outcome executable code is a concrete instance of the model. In this paper, we will focus on functional models for which the abstract description describes required functionality of the program, constraints on the hardware, and possible optimization criteria. At this level of description, there are many details that need to be filled. For example, each function may have many types of implementation, and each type can be implemented in itself in many different ways. An adder for example may be implemented as a simple ripple-carry adder, or as a Fourier transformed adder in which computation is done in quantum Fourier space. The result of the adder can be put out of place on a newly allocated register of qubits, leaving the two input registers intact, or can replace one or both of the input variables. Each operation within the adder may use any number of auxiliary qubits (’auxiliaries’) in order to perform the operation faster or with fewer two-qubit gates. Finally, the adder may be implemented to any level of accuracy, again affecting the number of qubits and two-qubit gates that the operation will consume. In all those cases, the functionality prescribed by the user remains intact [^1] Playing with the accuracy degree of freedom results in different functionalities of the adder, however we consider a likely model in which only the accuracy of the entire program is specified, and not the accuracy of each component.. Beyond the single function, unspecified details may also relate to the interaction between functions. For example, different functions may or may not share auxiliary or functional qubits, depending on the scope of the function and whether its results are required downstream, and also on the cost of disentangling and resetting used qubits. More generally, the decision on which qubits, if at all, to disentangle and reuse, as well as the ordering of functions using the same auxiliary qubits, is a very important ’detail’ from the implementation point of view, but is of no concern to the user from a functional model point of view. Such considerations of ’filling in the details’ between wide levels of abstraction are the basis of many electronic design automation (EDA) technologies, specifically synthesis from high-level functional and nonfunctional models to register-transfer-level (RTL) models and then to logic, netlist, and physical-level implementations. Given the circuit model of quantum computation, it should not be surprising that this general synthesis approach may also be applicable in the quantum domain. In this paper, we present the application of the EDA synthesis approach to quantum computing. Throughout the paper, we use the word synthesis as its meaning in the EDA industry, which is much more general than the term sometimes used in quantum computing [41] - meaning the decomposition of a specific unitary into a given gate set. Our results below demonstrate the viability of this approach for creating robust quantum software at scale. When developing the technology, we tried to preserve as much as possible all the guidelines and best practices of the conventional principles of EDA synthesis. The challenges we encountered in applying those principles to the nature of quantum programs were already significant enough so that we did not attempt to invent any new wheel when it was not necessary. Still, the peculiarities of quantum programs, including no fan-in or fan-out, no copy, disentanglement as means for freeing qubit resources, long-range correlations, inability to locally simulate sub-circuits, and many more – make the incorporation of these concepts into EDA synthesis principles non-trivial and challenging. In this paper, we mainly focus on the synthesis technology aspect per se. While we present modeling examples when needed, we do not delve into the research behind the definition of the quantum modeling constructs and the common quantum idioms that we found leading to those constructs. In addition, we do not describe here the challenges behind analysis and debugging of the automatically synthesized quantum program, and feedback into improving the model from this analysis. The modeling constructs and automatic program analysis are important parts of the EDA flow that we have again adapted into the quantum software cycle but are beyond the scope of this paper.
2 Results
We present results showing the power of our approach, in which a truly abstract description implies functional flexibility that brings about optimized gate-level quantum programs. We use Qmod [42] and the Classiq open library [43] to model the abstract functionality of the program and synthesize concrete programs under hardware constraints and optimization functions using the synthesis engine, as described in the Methods section 4 and the Supplementary Materials. We compare our results to those achieved by today’s state-of-the-art quantum coding tools. The problems we have chosen for demonstration are the most generic one can think of, and the improvement in the results compared to state-of-the-art tools is only due to the new principles introduced by our approach.
2.1 Quantum walk on a circle
Initially inspired by random classical walks, quantum walk [44] is an algorithmic technique that underlies many quantum algorithms, such as quantum search [45, 46] and Hamiltonian simulation based on Qubitization [47]. We consider the pedagogical example of a discrete quantum walk on a cyclic graph of $2^{N}$ nodes and a symmetric coin (see Ref. [48] for a comprehensive overview). The model includes two quantum variables, a qubit indicating the coin state and an array of qubits, $x$, representing the position of a walker along the circuit, i.e. an integer in $[0,2^{N}-1]$; see the schematic description in Fig. 1a.
The main quantum block of the algorithm is the increment function ($x\mathrel{+}=1\pmod{2^{N}}$). We focus on a specific, yet flexible design for this quantum function, given as a series of multi-controlled X (MCX) operations [^2] We note that there are various ways to model modular addition, e.g. by moving to Fourier space [49, 50]. However, we focus on the MCX based implementation as it highlights the rational of our approach. In fact, such an implementation was recently used to benchmark quantum hardware [51].. The MCX primitive has a variety of gate-level implementations, which differ in depth and number of auxiliaries [52, 53]. A high-level model (see Listings 1 and 2 in the Methods section) does not specify any such specific detail of the MCX implementation; this is determined dynamically by the synthesis engine given some global constraints for the full model. Panels (b) and (c) in Fig. 1 present two resulting quantum programs obtained by the synthesis engine for a circle with $2^{5}$ nodes, for two different optimization scenarios. We can see that when optimizing over the number of CX gates, the automated outcome includes two additional auxiliary qubits, allowing reduction of CX counts at the expense of the circuit’s width. Both auxiliaries are used for the first MCX operation, then only one of them is re-used for the two consecutive operations, and both auxiliaries are finally reused for the controlled decrement operations.
Figure 1: Quantum walk on a circular graph with $2^{N}$ nodes. (a) A quantum circuit diagram for a single walking step, defined by three operations: a Hadamard gate on the coin qubit, followed by increment and decrement on the position qubit array, controlled on the coin qubit being at state 0 and 1, respectively. (b, c): Two functionally-equivalent, different implementations of the model. In both examples the controlled-increment block is opened one hierarchy down, showing its explicit implementation. (b): The result of synthesizing the functional model with width optimization, giving a quantum circuit of 5 qubits and CX-counts of 274. (c): The result of synthesizing the functional model with CX-count optimization and constraining the width to 10 qubits. The synthesis outcome in this case is a quantum circuit of 7 qubits and CX-count of 120.
Trying to perform by hand the automation described in the last paragraph — decisions on local implementations out of multiple possibilities and auxiliary reuse management — is impractical. The effectiveness of this dynamic resource allocation becomes even more pronounced as we move to larger, more complex, and less structured models. In Fig. 2 we run the same quantum walk model for different values of circle sizes $2^{N}$, synthesizing with optimization over the CX count and constraining with two different values of maximal width. Our in-house library contains flexible MCX implementations that, given a sufficient number of auxiliaries, can reach a CX count of $O(n_{\rm ctrl})$ with $n_{\rm ctrl}$ being the size of the control variable, resulting in a scaling of $O(N^{2})$ of the quantum walk. From a software architecture perspective, any user can implement and provide any additional implementations of the MCX function (or any other function), and those will be taken into account by the synthesis engine just as the ones in our in-house library.
Fig. 2 also shows a comparison to the best-of-the-art model implementations using several other quantum software tools. In all those cases, implementations have a fixed CX-count for any number of control qubits in the function: exponential number in Qiskit [4] and Pennylane [5], and a cubic number in TKet [29]). We note that the libraries of these tools do have ad hoc, more efficient implementations for the MCX operation in terms of the CX gate count. However, when using these more efficient implementations, auxiliary qubits must be declared and treated as part of the function itself. Implementing the program with the compilers of these tools, which lack auxiliary awareness and management, results in a large circuit width because the auxiliaries of each function add up instead of being reused. In general terms, the underlying concept of all existing standard tools (including those that do include memory management [54, 35, 55]) is that they do not distinguish between the notion of a function and that of an implementation.
We finally note that one can go a further step and consider multiple implementations based on different designs for the increment function, which do not originate in the flexibility of the MCX building block, e.g., by introducing QFT-based or ripple-carry adders. For the sake of simplicity, we have not incorporated this degree of freedom in the simple quantum-walk experiment, and we consider this effect when dealing with larger problems in the next section.
Figure 2: Large scale quantum walk programs on a circle of $2^{N}$ nodes. The classiq curves are obtained by synthesizing with optimization over the CX counts and constraining the total width to 40 (blue dots) or 100 (brown stars). (a,b) The total CX counts, and the total generation time respectively (we stop at compilation times of $O(10^{3})$sec). The inset in panel (b) indicates the auxiliary usage. For comparison, the results for a similar implementation using various quantum platforms are plotted as well (see Methods for technical detail, and code for generating the results in Ref. [56]).
2.2 Quantum Singular Value Transform
We now examine a more applicative and less-structured example, applying a Quantum Singular Value Transform (QSVT) on a sparse block-encoded matrix. QSVT is a quantum primitive that is at the core of many algorithms such as quantum linear solvers, Hamiltonian simulation, and amplitude loading (see Ref. [57] and references therein). It supports a wide range of applications going from Finance [58], through Quantum Machine Learning [59] to Plasma Physics [60].
The implementation of QSVT is based on the block encoding of some matrix $A$ into a larger unitary matrix $U_{A}$, which can then be executed on quantum hardware. We block encode a simple $2^{N}\times 2^{N}$ matrix
$A=\begin{pmatrix}0&0&0&\dots&0&0\\ 0.5&0&0.5&\dots&0&0\\ 0&0.5&0&0.5&\dots&0\\ \vdots&0&\ddots&0&\dots&0\\ \vdots&0&\ddots&0.5&0&0.5\\ \end{pmatrix}.$ (1)
Quantum speedup that is based on the QSVT methodology assumes that the problem, e.g., the linear equation to be solved or the Hamiltonian to be evolved, is sparse. We find that synthesizing a rather minimal use-case as Eq. (1) already shows significant improvement compared to state of the art compiler tools working without the synthesis principles.
The block encoding of $A$ can be defined with quantum adders and reflection operations; see the schematic model in Fig. 3(a) and technical details in the Methods Section 4. In particular, the encoding unitary is of size $2^{N+3}$, having 3 block qubits [^3] In Fig. 3 we add another quantum variable $|s\rangle$ for making the problem even less structured, supporting, e.g., superposition of block-encoded matrices.. Our in-house library includes two different implementations of an in-place adder function on a quantum variable of size $M$: a QFT adder [61], that uses no extra auxiliary qubits and has an $O(M^{2})$ CX-counts, and a ripple-carry adder [62] that uses $M+1$ auxiliaries and has an $O(M)$ CX-counts. Fig. 3 shows the implementations provided by the synthesis engine for the block encoding model with two different optimization scenarios.
Refer to captionRefer to caption
Figure 3: Block encoding model for the matrix $A$ in Eq. (1). (a): A schematic description of the model, the $I-|0\rangle\langle 0|$ function is defined with single-qubit gates and an MCX operation, see Methods. (b,c): Two different implementations for $N=20$ obtained by the synthesis engine when optimizing over CX-counts and constraining the width to 45 (b) or 90 (c) qubits. Different adders are chosen (QFT or ripple carry, respectively), given the different constraints, resulting in CX-counts of 1480 and 842 for a 44 and 47 qubits circuit in panels (b) and (c), respectively.
Given a block encoding model, one can plug it into a QSVT routine to block encode a polynomial of degree $l$, $P_{l}(A)$ [57]. The model corresponds to applications of $U_{A}$, its inverse, and rotations about the block states; see Methods. We generate a random odd polynomial of degree 3 and apply it to different values of the matrix size, see Fig. 4. Our optimized results for CX counts, compared to state-of-the-art results from other software tools, originates from the flexibility of choosing efficient implementations of controlled operations and quantum adders.
Figure 4: QSVT for a polynomial of matrix $A$ in Eq. (1) for varying matrix sizes $2^{N}$. (a,b) Total CX counts and total generation time, where the inset of panel (b) presents the auxiliary counts. The Classiq curves are obtained by optimizing over the CX counts and constraining the total width to 50 (brown stars) or 100 (blue dots).
3 Conclusions
We have discussed the application of EDA synthesis methods to the field of quantum software and have shown clear results that support this approach as the robust, scalable and long-lasting way to approach quantum software development. One thing we have not covered in detail are the affect of different optimization functions and in particular fitness to a given hardware. For example, the graphs of Figs. 2,4 show results for which the synthesis engine optimizes for a minimal number of CX gates. They would look very different if the user decided to optimize for minimal width, providing much narrower circuits to the user. In addition, taking target hardware into account may have a crucial effect on the synthesis. For example, in fully connected hardware the synthesis engine may be more likely to implement an MCX gate with more CX gates than when the target is sparsely connected hardware, because in the latter case the cost of each CX gate is very large because of the need for SWAPS. We stress that such decisions on how to implement an MCX or other high-level functions cannot be taken by a transpiler which accepts a low-level implementation as input. Only true abstraction and true dynamic synthesis, as in this paper, can make such crucial decisions. Lastly, we stress that the decision on how to implement a function may be different for each instance of the function in the program, depending on the resources available locally for the function at this part of the program.
While we laid the foundations of the EDA-based synthesis concept and concretized it by an industry-grade software platform, there is still much to be done. A main item that is not yet fully implemented is the management of the accuracy resource. Specifically, automatic distribution of the global required algorithmic accuracy between the individual blocks in the program. Moreover, for noisy intermediate scale quantum (NISQ) target devices, the requirement is to decide on the accuracy of each block both from the overall required algorithmic accuracy and from the competition with the hardware noise. Further into the future are items related to the co-design and co-synthesis of the logical quantum program and the error-correction code it runs upon in fault-tolerant scenarios. We have applied EDA methods for determining the positioning of logical qubit patches in fault-tolerant schemes, such as to minimize two-qubit logical gate collisions and minimize overall program runtime. However, co-creating the logical program together with the error-correcting code is highly challenging as it involves the co-synthesis of coupled systems within two very different levels of abstraction.
Our technology adapts additional methods from EDA that were not described in this paper. These include the design of Qmod [42] as the high-level modeling language for quantum; debugging and analysis approaches to the quantum program coming out of the synthesis engine; modeling and synthesizing hybrid quantum-classical programs; visualization tools and graphical editors for quantum models and quantum programs; and nonfunctional modeling idioms for quantum. Examples of such idioms are requirements for levels of entanglement in specific parts of the program, phase difference between different parts of the program, and more. In all those areas, as well as in the main topic presented in this paper — synthesis of quantum programs from high-level functional models — there remains a huge amount of work to do in order to reach the level of a robust, scalable, reliable, usable, and fast software ready for addressing the demanding needs of quantum algorithm developers in a post-quantum-advantage industrial world.
4 Methods
A detailed description of the synthesis engine is provided in the Supplementary Materials. Here we provide some technical details needed to perform the experiments of Sec. 2, while the corresponding code is available in Ref. [56].
4.1 Numerical study details
The results presented in the paper are obtained by modeling simple quantum models on various quantum platforms, using their Python SDK package for modeling and compilation per platform. The software versions used in this paper are: Classiq 0.60.0, Qiskit 1.2.4, PyTket 1.34.0, and PennyLane 0.39.0 with PennyLane-Catalyst compiler 0.9.0. For each platform, we use its quantum control operation, MCX function, basic single-qubit gate operations, and built-in quantum arithmetic operations (when available, otherwise we define it ourselves). In all cases, experts knowledgeable about the software tools made their best effort to apply the tools in the most reasonable way to achieve the task in question. The specific inputs to the tools are provided in Ref. [56].
In all platforms, we use the tool’s compilation and transpilation abilities to get a final quantum program with only one-qubit gate operations and CX operations. This allows a fair comparison for the CX counts per example. All runtime experiments were performed on a single Apple M1 pro, all on the same computer with the same software installed and the same environmental conditions. During run-time, the software in question was the only significant application running on the computer.
4.2 Quantum walk model
We discuss in detail the quantum model presented in Sect. 2.1. The model definition is accompanied with a Qmod code for highlighting the high-level description. A discrete quantum walk refers to the repetitions over ”coin flipping” followed by a conditional step according to the coin state. For the circular graph, the coin operation is defined by a a Hadamard gate, represented by the $2\times 2$ matrix $H=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$. This is a standard 1-qubit gate, available in all software tools and languages. A clockwise (counter clockwise) step is given by an increment (decrement) operation, referring to an in-place addition (subtraction) by 1, modulo the number of nodes $2^{N}$. We design the increment function as a cascade of MCX operations: {mdframed}[backgroundcolor=light-gray, roundcorner=10pt,leftmargin=1, rightmargin=1, innerleftmargin=15, innertopmargin=15,innerbottommargin=15, outerlinewidth=1, linecolor=light-gray]
Listing 1: An implemetation of an increment operation
qfunc my_mcx(x: qnum, y: qbit) {
control (x==((2**x.size)-1)) {
X(y);
qfunc increment(x: qbit[]) {
repeat (i: x.len - 1) {
my_mcx(x[0:(x.len - 1) - i],
x[(x.len - 1) - i]);
X(x[0]);
Then, a single walking step is defined via control and invert operations: {mdframed}[backgroundcolor=light-gray, roundcorner=10pt,leftmargin=1, rightmargin=1, innerleftmargin=15, innertopmargin=15,innerbottommargin=15, outerlinewidth=1, linecolor=light-gray]
Listing 2: An implemetation of a quantum walk on a circle
qfunc single_step(coin: qbit, x: qnum)
H(coin);
control (coin == 0) {
increment(x);
control (coin == 1) {
invert {
increment(x);
The models in all other platforms are designed in the same fashion, using the platforms’ building blocks such as quantum control operation, inversion, and MCX gates [56].
4.3 QSVT on the block-encoded $A$
We block encode the matrix $A$ in Eq. (1) as a product of $A_{0}=0.5\sum^{2^{N}-2}_{i=0}(|i\rangle\langle i+1|+0.5|2^{N}-1\rangle\langle 0|+c.c$ and $A_{1}=I-|0\rangle\langle 0|$, , and apply block-encoding multiplication. The block encoding of $A_{0}$ is done with 2 block qubits, according to the procedure described in [63] for Toeplitz matrices [see Eq. (55) therein]. The matrix $A_{1}$ can be block encoded as a Linear Combination of Unitaries (LCU) for the unitaries $I$ and $I-2|0\rangle\langle 0|$. The latter unitary is the reflection around zero operation appearing in Grover’s Search algorithm. The Qmod code for block encoding $A$ is given in Listing 3.
{mdframed}
[backgroundcolor=light-gray, roundcorner=10pt,leftmargin=1, rightmargin=1, innerleftmargin=15, innertopmargin=15,innerbottommargin=15, outerlinewidth=1, linecolor=light-gray]
Listing 3: An implemetation of a block encoding A
qfunc be_amat0(data: qbit[], block: qbit[]) {
del_qubit: qbit;
select: qbit;
packed: qnum<data.size + 1>;
within {
block -> {select, del_qubit};
{data, del_qubit} -> packed;
hadamard_transform(select);
} apply {
control (select == 0) {
packed += 2;
packed += -1;
qfunc be_projection(x: qbit[], aux: qbit) {
within {
H(aux);
} apply {
control (aux == 0) {
reflect_about_zero(x);
qfunc be_amat(data: qbit[], block: qbit[]) {
be_amat0(data, block[0:2]);
be_projection(data, block[2]);
The QSVT scheme for a polynomial with a given parity is a standard quantum primitive. It involves iterations over application of the block encoding unitary, its inverse, one-qubit rotations, and reflection operations (see Fig. 1 in Ref. [59]). We use the classiq-library [43] functions to obtain the QSVT application, as well as for the reflection about zero unitary. The QSVT function in Classiq open library accepts a projection operator, given in Listing 4.
{mdframed}
[backgroundcolor=light-gray, roundcorner=10pt,leftmargin=1, rightmargin=1, innerleftmargin=15, innertopmargin=15,innerbottommargin=15, outerlinewidth=1, linecolor=light-gray]
Listing 4: An implemetation of a reflection in the block space
qfunc my_projector_controlled_phase(phase: real, block: qnum, aux: qbit) {
control (block == 0) {
X(aux);
RZ(phase, aux);
control (block == 0) {
X(aux);
[…]
Supplementary information
The article is accompanied by a Supplemental Materials describing the synthesis engine. The code for obtaining using other quantum softwares and compilers is available in a public repository [56].
Declarations
Author contribution
Classiq team developed, designed, and implemented the Synthesis engine, the Qmod language, and the classiq-library, and reviewed and adjusted the manuscript. T.G and Y.N, wrote the manuscript and I.R the Supplementary Materials. T.G ran the numerical experiments on the Synthesis engine and other quantum platforms.